OpenStudy (anonymous):

Emily wants to hang a painting in a gallery but the painting and frame must have an area of 31 square feet. The painting is 5 feet wide by 6 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined?

OpenStudy (anonymous):

L=5+x W=6+x A=L*W=(x+5)(x+6) expand away.